Nouabalé Ndoki (10.01.2012 to 12.31.2012)

You are here

PDF versionPDF version
October - December
Protocol Activities: 

CL protocol:
We’ve proceeded to our first temperature/humidity sensor calibration.
All the CL data are uploaded on the TEAM portal in a regular basis.

VG protocol:
3 permanent plots have been surveyed in December 2012.
987 stems (including 9 new recruits) have been measured and recorded

CT protocol:
30 cameras have been successfully deployed during the two-week field mission.
29 cameras traps were placed at their usual ID points.
Camera CT-NNN-227 had to be moved to a nearby and geo-referenced location as its initial host tree died.

Schedule Problems: 

Our CT and VG fiield activities were initially scheduled for end November. We delayed the starting sampling period due to heavy frequent rains.